OpenStudy (anonymous):

Simplify (-8 + 8i) - (5 + 4i) -17 -9 -13 + 12i -13 + 4i

OpenStudy (anonymous):

=-8+8i-5-4i=(-8-5)+(8-4)i=-13+4i option d